# Optimization Insights into Deep Diagonal Linear Networks

### v2

@@ 3 Conclusion and Perspectives
In this work, we have studied Deep Diagonal Linear Networks and broaden our understanding of their training. In particular, we have analyzed the trajectories of the original variable when a Gradient Flow is applied to its layers. In this way, we were able to extract a sufficient initialization condition to prove theoretical results both on the optimization and the bias of this model. We show indeed that such a training induces a Mirror Flow on the original variable and ensures a linear decrease of the error on the loss function as long as it satisfies the Polyak-Łojasiewicz inequality. This study also reveals the negative impact of initializing the network with small values in terms of convergence speed.
Despite its apparent simplicity, this model is still tricky to study and it remains as an open question whether it is possible to compute explicitly the entropy function associated to the Mirror Flow. It is then a natural question remaining that would also inform us on the implicit bias underlying. It is also noteworthy that the linearity of this model may prevent us from fully understanding the generalization properties of real-world neural networks. This highlights the potential necessity for incorporating non-linearity in future works. From an optimization perspective, further research could be conducted to investigate the influence of stochasticity, momentum, and more generally, the optimization method applied during training.
@@ A.1 Background: A direct strategy to analyze Diagonal Linear Networks
In this section, we briefly state the strategy proposed by [Woodworth et al., 2020] to analyze the training of a Diagonal Linear Network through Gradient Flow. The proposed approach to derive an implicit bias benefits from its simplicity and is general enough to be extended to other simple models. To emphasize this point, we apply it also to Deep Redundant Linear Network in Appendix A.3.
@@
[…]
@@
[…]
@@
In order to achieve this, [Woodworth et al., 2020] notice that $\theta_{\infty}=\Psi(X^{T}\nu)$ (where $\Psi$ is defined in equation 17) for some $\nu\in\mathbb{R}^{n}$. Indeed,
[…]
@@
$\left\{\begin{gathered}X\theta_{\infty}=y\\ \exists\omega\in\mathbb{R}^{n},\leavevmode\nobreak\ \nabla\mathcal{Q}(\theta_{\infty})=X^{T}\omega\end{gathered}\right.$ (19)
@@ A.2 Detailed computations for Diagonal Linear Networks
[…]
It is then trivial to observe that $(z_{+},z_{-})$ defined as $(u+v,u-v)$ is a solution of:
[…]
@@
[…]
@@ A.3 Study of a Deep Redundant Linear Network
In this section, we emphasize that the strategy stated in Section A.1 is general enough to be applied to other similar parameterizations. We briefly analyze the implicit features of a Deep Redundant Linear Network defined through similar computations. The stated results can be seen as a simplified version of that obtained in [Woodworth et al., 2020, Moroshko et al., 2020, Chou et al., 2023a] in the $L$-layer case.
We consider the parameterization $\theta=u^{\odot L}$ where the number of layers $L\in\mathbb{N}$ is strictly greater than $2$. We also suppose that the loss function $\mathcal{L}$ can be written as $\mathcal{L}(\theta)=\ell(X\theta-y)$ for any $\theta\in\mathbb{R}^{d}$, with $\ell:\mathbb{R}^{n}\rightarrow\mathbb{R}$ which reaches its minimum at 00.
@@
Supposing that for any $i\in\mathbb{R}^{d}$, $u_{i}(t)$ is different from 00 for any $t\geqslant 0$ (which can be done by enforcing positivity of $X$, $y$ and the initialization), we can write that
@@
[…]
@@ A.4 Detailed proof of Theorem 1
=$\mathcal{M}=\left\{w\in\mathbb{R}^{L\times d},\leavevmode\nobreak\ \forall i\in[d],\leavevmode\nobreak\ \#\left\{j\in[L],\leavevmode\nobreak\ w_{(i-1)L+(j-1)}=0\right\}\leqslant 1\right\}.$ (33)
@@
$\mathcal{M}=\bigcap_{i\in[d]}\left\{w\in\mathbb{R}^{L\times d},\leavevmode\nobreak\ \#\left\{j\in[L],\leavevmode\nobreak\ w_{(i-1)L+(j-1)}=0\right\}\leqslant 1\right\}.$
@@
$\mathcal{M}^{C}$ $=\bigcup_{i\in[d]}\left\{w\in\mathbb{R}^{L\times d},\leavevmode\nobreak\ \exists(j,k)\in[L]^{2},\leavevmode\nobreak\ j\neq k,\leavevmode\nobreak\ w_{(i-1)L+(j-1)}=w_{(i-1)L+(k-1)}=0\right\}$
$=\bigcup_{i\in[d]}\bigcup_{(j,k)\in[L]^{2},\leavevmode\nobreak\ j\neq k}\left\{w\in\mathbb{R}^{L\times d},\leavevmode\nobreak\ w_{(i-1)L+(j-1)}=w_{(i-1)L+(k-1)}=0\right\}.$
@@
$\forall j\in[Ld],\leavevmode\nobreak\ \left(\nabla G_{i}(w)\right)_{j}=\left\{\begin{aligned} \prod_{k=(i-1)L+1,\leavevmode\nobreak\ k\neq j}^{iL}&w_{k}\quad\text{if}\leavevmode\nobreak\ j\in\llbracket(i-1)L+1,iL\rrbracket,\\ &0\quad\text{else.}\end{aligned}\right.$ (34)
@@
$\cdot$ We finally need to check that [Li et al., 2022, Assumption 3.5] is satisfied by $G$, i.e that the domain of the flows induced by its gradient vector fields is pairwise symmetric. We refer the curious reader to the reference for more details, but we simply notice that since $\nabla G_{j}(w)$ and $\nabla G_{k}(w)$ have non zero entries on disjoint sets of components (if $j\neq k$), this assumption is directly satisfied.
From this point, we apply [Li et al., 2022, Theorem 4.8] and conclude.

### v3

@@ 3 Conclusion and Perspectives
In this work, we analyzed the optimization dynamics of Deep Diagonal Linear Networks by studying the trajectory of the effective parameter induced by gradient flow on the layers weights variables. This approach allows to identify mild initialization conditions under which the dynamics in the effective parameter space follow a mirror-flow structure and to derive convergence guarantees, including linear convergence under a Polyak–Łojasiewicz condition. Our analysis highlights the detrimental effect of excessively small initializations on the convergence speed. Although the model we consider corresponds to a simple learning architecture, its dynamics and analysis remain surprisingly intricate. An open question is whether the entropy function associated with the induced mirror flow can be characterized explicitly. This would further clarify the underlying implicit geometry as well as the induced bias. Clearly considering non linear models, e.g. homogenous nonlinearities, would be a step towards more practical learning architectures Finally, we note that from an optimization standpoint, incorporating stochasticity, momentum, or alternative update rules offers additional avenues for understanding how algorithmic choices interact with overparameterized parametrizations.
@@ A.1 Background: A direct strategy to analyze Diagonal Linear Networks
In this section, we briefly state the strategy proposed by [37] to analyze the training of a Diagonal Linear Network through Gradient Flow. The proposed approach to derive an implicit bias benefits from its simplicity and is general enough to be extended to other simple models. To emphasize this point, we apply it also to Deep Redundant Linear Network in Appendix A.3.
@@
$\left\{\begin{aligned} &\dot{u}(t)+v(t)\odot\nabla\mathcal{L}(\theta(t))=0\\ &\dot{v}(t)+u(t)\odot\nabla\mathcal{L}(\theta(t))=0.\end{aligned}\right$. (15)
@@
where $\xi:t\mapsto-\int_{0}^{t}\nabla\mathcal{L}(\theta(s))ds$ is the solution of
@@
In order to achieve this, [37] notice that $\theta_{\infty}=\Psi(X^{T}\nu)$ (where $\Psi$ is defined in equation 16) for some $\nu\in\mathbb{R}^{n}$. Indeed,
$\lim\limits_{t\rightarrow+\infty}\xi(t)=-\int_{0}^{+\infty}\nabla\mathcal{L}(\theta(s))ds=-X^{T}\underbrace{\int_{0}^{+\infty}\nabla\ell(X\theta(s)-y)ds}_{:=-\nu}.$
@@
$\left\{\begin{gathered}X\theta_{\infty}=y\\ \exists\omega\in\mathbb{R}^{n},\penalty 10000\ \nabla\mathcal{Q}(\theta_{\infty})=X^{T}\omega\end{gathered}\right$. (18)
@@ A.2 Detailed computations for Diagonal Linear Networks
$\left\{\begin{aligned} &\dot{u}(t)+v(t)\odot\nabla\mathcal{L}(\theta(t))=0\\ &\dot{v}(t)+u(t)\odot\nabla\mathcal{L}(\theta(t))=0.\end{aligned}\right$. (22)
It is then trivial to observe that $(z_{+},z_{-})$ defined as (u+v,uv)(u+v,u-v) is a solution of:
$\left\{\begin{aligned} &\dot{z}_{+}(t)+z_{+}(t)\odot\nabla\mathcal{L}(\theta(t))=0\\ &\dot{z}_{-}(t)-z_{-}(t)\odot\nabla\mathcal{L}(\theta(t))=0\end{aligned}\right$., (23)
@@
$\left\{\begin{aligned} &z_{+}(t)=z_{+}(0)e^{\xi(t)}\\ \ &z_{-}(t)=z_{-}(0)e^{-\xi(t)}\end{aligned}\right$., (24)
where $\xi:t\mapsto-\int_{0}^{t}\nabla\mathcal{L}(\theta(s))ds$. Note by defining $\xi$ in this way, it holds that
@@ A.3 Study of a Deep Redundant Linear Network
In this section, we emphasize that the strategy stated in Section A.1 is general enough to be applied to other similar parameterizations. We briefly analyze the implicit features of a Deep Redundant Linear Network defined through similar computations. The stated results can be seen as a simplified version of that obtained in [37, 26, 10] in the $L$-layer case.
We consider the parameterization $\theta=u^{\odot L}$ where the number of layers $L\in\mathbb{N}$ is strictly greater than $2$. We also suppose that the loss function $\mathcal{L}$ can be written as $\mathcal{L}(\theta)=\ell(X\theta-y)$ for any $\theta\in\mathbb{R}^{d}$, with $\ell:\mathbb{R}^{n}\rightarrow\mathbb{R}$ which reaches its minimum at 0.
@@
Supposing that for any $i\in\mathbb{R}^{d}$, $u_{i}(t)$ is different from 0 for any $t\geqslant 0$ (which can be done by enforcing positivity of $X$, $y$ and the initialization), we can write that
@@
where $\xi:t\mapsto-\int_{0}^{t}\nabla\mathcal{L}(\theta(s))ds$. The original variable $\theta$ then satisfies:
@@ A.4 Detailed proof of Theorem 1
$\mathcal{M}=\left\{w\in\mathbb{R}^{L\times d},\penalty 10000\ \forall i\in[d],\penalty 10000\ \#\left\{j\in[L],\penalty 10000\ w_{(i-1)L+(j-1)}=0\right\}\leqslant 1\right\}.$ (32)
@@
$\mathcal{M}=\bigcap_{i\in[d]}\left\{w\in\mathbb{R}^{L\times d},\penalty 10000\ \#\left\{j\in[L],\penalty 10000\ w_{(i-1)L+(j-1)}=0\right\}\leqslant 1\right\}.$
@@
$\mathcal{M}^{C}$ $=\bigcup_{i\in[d]}\left\{w\in\mathbb{R}^{L\times d},\penalty 10000\ \exists(j,k)\in[L]^{2},\penalty 10000\ j\neq k,\penalty 10000\ w_{(i-1)L+(j-1)}=w_{(i-1)L+(k-1)}=0\right\}$
$=\bigcup_{i\in[d]}\bigcup_{(j,k)\in[L]^{2},\penalty 10000\ j\neq k}\left\{w\in\mathbb{R}^{L\times d},\penalty 10000\ w_{(i-1)L+(j-1)}=w_{(i-1)L+(k-1)}=0\right\}.$
@@
$\forall j\in[Ld],\penalty 10000\ \left(\nabla G_{i}(w)\right)_{j}=\left\{\begin{aligned} \prod_{k=(i-1)L+1,\penalty 10000\ k\neq j}^{iL}&w_{k}\quad\text{if}\penalty 10000\ j\in\llbracket(i-1)L+1,iL\rrbracket,\\ &0\quad\text{else.}\end{aligned}\right$. (33)
@@
$\cdot$ We finally need to check that [21, Assumption 3.5] is satisfied by $G$, i.e that the domain of the flows induced by its gradient vector fields is pairwise symmetric. We refer the curious reader to the reference for more details, but we simply notice that since $\nabla G_{j}(w)$ and $\nabla G_{k}(w)$ have non zero entries on disjoint sets of components (if $j\neq k$), this assumption is directly satisfied.
From this point, we apply [21, Lemma 4.8 and Theorem 4.9] and conclude the first part. To prove equation 8, note that
$\frac{d}{dt}\mathcal{L}(\theta(t))$ $=\langle\nabla\mathcal{L}(\theta(t)),\dot{\theta}(t)\rangle$
$=-\langle\nabla^{2}Q(\theta(t))\dot{\theta}(t),\dot{\theta}(t)\rangle$
$\leq 0,$
which implies that $\mathcal{L}(\theta(t))$ is non increasing. Moreover,
$\frac{d}{dt}D_{Q}(\theta^{*},\theta)$ $=\frac{d}{dt}\Big[Q(\theta^{*})-Q(\theta(t))-\langle\nabla Q(\theta(t)),\theta^{*}-\theta(t)\rangle\Big]$
$=-\langle\nabla Q(\theta(t)),\dot{\theta}\rangle+\langle\nabla Q(\theta(t)),\dot{\theta}\rangle-\langle\frac{d}{dt}\nabla Q(\theta(t)),\theta^{*}-\theta(t)\rangle$
$=\langle\nabla\mathcal{L}(\theta(t)),\theta^{*}-\theta(t)\rangle$
$\leq-(\mathcal{L}(\theta(t))-\mathcal{L}^{*}).$
Therefore, $(\mathcal{L}(\theta(t))-\mathcal{L}^{*})$ is decreasing and integrable, and
$D_{Q}(\theta^{*},\theta(t))+\int_{0}^{t}(\mathcal{L}(\theta(s))-\mathcal{L}^{*})ds\leq D_{Q}(\theta^{*},\theta(0)).$
The statement then follows.